# Model Context Protocol (MCP): Landscape, Security Threats, and Future Research Directions

### v2

@@ Abstract.
The Model Context Protocol (MCP) is a standardized interface designed to enable seamless interaction between AI models and external tools and resources, breaking down data silos and facilitating interoperability across diverse systems. This paper provides a comprehensive overview of MCP, focusing on its core components, workflow, and the lifecycle of MCP servers, which consists of three key phases: creation, operation, and update. We analyze the security and privacy risks associated with each phase and propose strategies to mitigate potential threats. The paper also examines the current MCP landscape, including its adoption by industry leaders and various use cases, as well as the tools and platforms supporting its integration. We explore future directions for MCP, highlighting the challenges and opportunities that will influence its adoption and evolution within the broader AI ecosystem. Finally, we offer recommendations for MCP stakeholders to ensure its secure and sustainable development as the AI landscape continues to evolve.
@@ 1. Introduction
In recent years, the vision of autonomous AI agents capable of interacting with a wide range of tools and data sources has gained significant momentum. This progress accelerated in 2023 with the introduction of function calling by OpenAI, which allowed language models to invoke external APIs in a structured way (OpenAI, 2023b). This advancement expanded the capabilities of LLMs, enabling them to retrieve real-time data, perform computations, and interact with external systems. As function calling gained adoption, an ecosystem formed around it. OpenAI introduced the ChatGPT plugin (OpenAI, 2023a), allowing developers to build callable tools for ChatGPT. LLM app stores such as Coze (ByteDance, 2024) and Yuanqi (Tencent, 2024) have launched their plugin stores, supporting tools specifically designed for their platforms. Frameworks like LangChain (LangChain, 2022) and LlamaIndex (Liu, 2022) provided standardized tool interfaces, making it easier to integrate LLMs with external services. Other AI providers, including Anthropic, Google, and Meta, introduced similar mechanisms, further driving adoption. Despite these advancements, integrating tools remains fragmented. Developers must manually define interfaces, manage authentication, and handle execution logic for each service. Function calling mechanisms vary across platforms, requiring redundant implementations. Additionally, current approaches rely on predefined workflows, limiting AI agents’ flexibility in dynamically discovering and orchestrating tools.
In late 2024, Anthropic introduced the Model Context Protocol (MCP)(Anthropic, 2024b), a general-purpose protocol standardizing AI-tool interactions. Inspired by the Language Server Protocol (LSP) (Gunasinghe and Marcus, 2021), MCP provides a flexible framework for AI applications to communicate with external tools dynamically. Instead of relying on predefined tool mappings, MCP allows AI agents to autonomously discover, select, and orchestrate tools based on task context. It also supports human-in-the-loop mechanisms, enabling users to inject data or approve actions as needed. By unifying interfaces, MCP simplifies the development of AI applications and improves their flexibility in handling complex workflows. Since its release, MCP has rapidly grown from a niche protocol to a key foundation for AI-native application development. A thriving ecosystem has emerged, with thousands of community-driven MCP servers enabling model access to systems like GitHub (Protocol, 2024a), Slack (Protocol, 2024b), and even 3D design tools like Blender (ahujasid, 2025). Tools like Cursor (Cursor, 2025) and Claude Desktop (Anthropic, 2024a) demonstrate how MCP clients can extend their capabilities by installing new servers, turning developer tools, productivity platforms, and creative environments alike into multi-modal AI agents.
@@
Despite the rapid adoption of MCP, its ecosystem is still in the early stages, with key areas such as security, tool discoverability, and remote deployment lacking comprehensive solutions. These issues present untapped opportunities for further research and development. Although MCP is widely recognized for its potential in the industry, it has not yet been extensively analyzed in academic research. This gap in research motivates this paper, which provides the first analysis of the MCP ecosystem, examining its architecture and workflow, defining the lifecycle of MCP servers, and identifying potential security risks at each stage, such as installer spoofing and tool name conflict. Through this study, we present a thorough exploration of MCP’s current landscape and offer a forward-looking vision that highlights key implications, outlines future research directions, and addresses the challenges that must be overcome to ensure its sustainable growth.
@@
  2. (2)
    We identify the key components of MCP servers and define their lifecycle, encompassing the stages of creation, operation, and update. We also highlight potential security risks associated with each phase, offering insights into safeguarding AI-to-tool interactions.
  3. (3)
@@
The remainder of this paper is structured as follows: § 2 compares tool invocation with and without MCP, highlighting the motivation for this study. § 3 outlines the architecture of MCP, detailing the roles of the MCP host, client, and server, as well as the lifecycle of the MCP server. § 4 examines the current MCP landscape, focusing on key industry players and adoption trends. § 5 analyzes security and privacy risks across the MCP server lifecycle and proposes mitigation strategies. § 6 explores implications, future challenges, and recommendations to enhance MCP’s scalability and security in dynamic AI environments. § 7 reviews prior work on tool integration and security in LLM applications. Finally, § 8 concludes the whole paper.
@@ 2.1. AI Tooling
Figure 1. Tool invocation with and without MCP.
@@ 2.1.2. Standardized Plugin Interfaces
To reduce the complexity of manual wiring, plugin-based interfaces such as OpenAI ChatGPT Plugins, introduced in November 2023 (OpenAI, 2023a), allowed AI models to connect with external tools through standardized API schemas like OpenAPI. For example, in the OpenAI Plugin ecosystem, plugins like Zapier allowed models to perform predefined actions, such as sending emails or updating CRM records. However, these interactions were often one-directional and could not maintain state or coordinate multiple steps in a task. New LLM app stores (Zhao et al., 2024) such as ByteDance Coze (ByteDance, 2024) and Tencent Yuanqi (Tencent, 2024) have also emerged, offering a plugin store for web services. While these platforms expanded available tool options, they created isolated ecosystems where plugins are platform-specific, limiting cross-platform compatibility and requiring duplicate maintenance efforts. MCP stands out by being open-source and platform-agnostic, enabling AI applications to engage in rich two-way interactions with external tools, facilitating complex workflows.
@@ 2.1.3. AI Agent Tool Integration
The emergence of AI agent frameworks like LangChain (LangChain, 2022) and similar tool orchestration frameworks provided a structured way for models to invoke external tools through predefined interfaces, improving automation and adaptability (Xi et al., 2023). However, integrating and maintaining these tools remained largely manual, requiring custom implementations and increasing complexity as the number of tools grew. MCP simplifies this process by offering a standardized protocol that enables AI agents to seamlessly invoke, interact with, and chain multiple tools through a unified interface. This reduces manual configuration and enhances task flexibility, allowing agents to perform complex operations without extensive custom integration.
@@
Contextual information retrieval methods, such as RAG, leverage vector-based search to retrieve relevant knowledge from databases or knowledge bases, enabling models to supplement responses with up-to-date information (Fan et al., 2024; Cuconasu et al., 2024). While this approach addressed the problem of knowledge cutoff and improved model accuracy, it was limited to passive retrieval of information. It did not inherently allow models to perform active operations, such as modifying data or triggering workflows. For example, a RAG-based system could retrieve relevant sections from a product documentation database to assist a customer support AI. However, if the AI needed to update customer records or escalate an issue to human support, it could not take action beyond providing textual responses. MCP extends beyond passive information retrieval by enabling AI models to interact with external data sources and tools actively, facilitating both retrieval and action in a unified workflow.
@@ 2.2. Motivation
MCP has rapidly gained traction in the AI community due to its ability to standardize how AI models interact with external tools, fetch data, and execute operations. By addressing the limitations of manual API wiring, plugin interfaces, and agent frameworks, MCP has the potential to redefine AI-to-tool interactions and enable more autonomous and intelligent agent workflows. Despite its growing adoption and promising potential, MCP is still in its early stages, with an evolving ecosystem that remains incomplete. Many key aspects, such as security and tool discoverability, are yet to be fully addressed, leaving ample room for future research and improvement. Moreover, while MCP has gained rapid adoption in the industry, it is still largely unexplored in academia.
@@ 3.1. Core Components
Figure 2. The workflow of MCP.
@@ 3.1.2. MCP Client
The MCP client acts as an intermediary within the host environment, managing communication between the MCP host and one or more MCP servers. It initiates requests to MCP servers, queries available functions, and retrieves responses that describe the server’s capabilities. This ensures seamless interaction between the host and external tools. In addition to managing requests and responses, the MCP client processes notifications from MCP servers, providing real-time updates about task progress and system status. It also performs sampling to gather data on tool usage and performance, enabling optimization and informed decision-making. The MCP client communicates through the transport layer with MCP servers, facilitating secure, reliable data exchange and smooth interaction between the host and external resources.
@@ 3.1.3. MCP Server
  • Tools: Enabling external operations. Tools allow the MCP server to invoke external services and APIs to execute operations on behalf of AI models. When the client requests an operation, the MCP server identifies the appropriate tool, interacts with the service, and returns the result. For instance, if an AI model requires real-time weather data or sentiment analysis, the MCP server connects to the relevant API, retrieves the data, and delivers it to the host. Unlike traditional function calling, which requires multiple steps and separates invocation from execution, Tools of MCP servers streamline this process by allowing the model to autonomously select and invoke the appropriate tool based on context. Once configured, these tools follow a standardized supply-and-consume model, making them modular, reusable, and easily accessible to other applications, enhancing system efficiency and flexibility.
@@ 3.3. MCP Server Lifecycle
The MCP server lifecycle as shown in Figure 3 consists of three key phases: creation, operation, and update. Each phase defines critical activities that ensure the secure and efficient functioning of the MCP server, enabling seamless interaction between AI models and external tools, resources, and prompts.
Figure 3. MCP servers components and lifecycle.
@@ 3.3.1. MCP Server Components
The MCP server is responsible for managing external tools, data sources, and workflows, providing AI models with the necessary resources to perform tasks efficiently and securely. It comprises several key components that ensure smooth and effective operations. Metadata includes essential information about the server, such as its name, version, and description, allowing clients to identify and interact with the appropriate server. Configuration involves the source code, configuration files, and manifest, which define the server’s operational parameters, environment settings, and security policies. Tool list stores a catalog of available tools, detailing their functionalities, input-output formats, and access permissions, ensuring proper tool management and security. Resources list governs access to external data sources, including web APIs, databases, and local files, specifying allowed endpoints and their associated permissions. Finally, Prompts and Templates include pre-configured task templates and workflows that enhance the efficiency of AI models in executing complex operations. Together, these components enable MCP servers to provide seamless tool integration, data retrieval, and task orchestration for AI-powered applications.
@@ 3.3.2. Creation Phase
The creation phase is the initial stage of the MCP server lifecycle, where the server is registered, configured, and prepared for operation. This phase involves three key steps. Server registration assigns a unique name and identity to the MCP server, allowing clients to discover and connect to the appropriate server instance. Installer deployment involves installing the MCP server and its associated components, ensuring that the correct configuration files, source code, and manifests are in place. Code integrity verification validates the integrity of the server’s codebase to prevent unauthorized modifications or tampering before the server becomes operational. Successful completion of the creation phase ensures that the MCP server is ready to handle requests and interact securely with external tools and data sources.

### v3

@@ Abstract.
The Model Context Protocol (MCP) is an emerging open standard that defines a unified, bi-directional communication and dynamic discovery protocol between AI models and external tools or resources, aiming to enhance interoperability and reduce fragmentation across diverse systems. This paper presents a systematic study of MCP from both architectural and security perspectives. We first define the full lifecycle of an MCP server, comprising four phases (creation, deployment, operation, and maintenance), further decomposed into 16 key activities that capture its functional evolution. Building on this lifecycle analysis, we construct a comprehensive threat taxonomy that categorizes security and privacy risks across four major attacker types: malicious developers, external attackers, malicious users, and security flaws, encompassing 16 distinct threat scenarios. To validate these risks, we develop and analyze real-world case studies that demonstrate concrete attack surfaces and vulnerability manifestations within MCP implementations. Based on these findings, the paper proposes a set of fine-grained, actionable security safeguards tailored to each lifecycle phase and threat category, offering practical guidance for secure MCP adoption. We also analyze the current MCP landscape, covering industry adoption, integration patterns, and supporting tools, to identify its technological strengths as well as existing limitations that constrain broader deployment. Finally, we outline future research and development directions aimed at strengthening MCP’s standardization, trust boundaries, and sustainable growth within the evolving ecosystem of tool-augmented AI systems. All collected data and implementation examples are publicly available at https://github.com/security-pride/MCP_Landscape.
@@ 1. Introduction
In recent years, the vision of autonomous AI agents capable of interacting with a wide range of tools and data sources has gained significant momentum. This progress accelerated in 2023 with the introduction of function calling by OpenAI, which allowed language models to invoke external APIs in a structured way (OpenAI, 2023b). This advancement expanded the capabilities of LLMs, enabling them to retrieve real-time data, perform computations, and interact with external systems. As function calling gained adoption, an ecosystem formed around it. OpenAI introduced the ChatGPT plugin (OpenAI, 2023a), allowing developers to build callable tools for ChatGPT. LLM app stores such as Coze (ByteDance, 2024) and Yuanqi (Tencent, 2024) have launched their plugin stores, supporting tools specifically designed for their platforms. Frameworks like LangChain (LangChain, 2022) and LlamaIndex (Liu, 2022) provided standardized tool interfaces, making it easier to integrate LLMs with external services. Other AI providers, including Anthropic, Google, and Meta, introduced similar mechanisms, further driving adoption. Despite these advancements, integrating tools remains fragmented. Developers must manually define interfaces, manage authentication, and handle execution logic for each service. Function calling mechanisms vary across platforms, requiring redundant implementations. Moreover, existing agent frameworks already support a degree of autonomous tool selection, but they generally operate over predefined or hardcoded integrations, limiting interoperability and long-term maintainability.
In late 2024, Anthropic launched the Model Context Protocol (MCP) (Anthropic, 2024b), a universal protocol for standardizing the definition, discovery, and invocation of external tools for AI applications. Drawing inspiration from the Language Server Protocol (LSP) (Gunasinghe and Marcus, 2021), MCP introduces several innovations that extend beyond conventional function calling. First, it provides a protocol-based standard that decouples tool implementation from usage, enabling developers to publish and describe external functions dynamically, independent of any single model or agent framework. Second, MCP supports dynamic discovery and schema negotiation: the client can list available tools at runtime, retrieve their capabilities, and invoke them in a uniform manner, without requiring prior hardcoding or platform-specific adapters. Third, MCP enables bi-directional communication channels, allowing not only model-to-tool requests but also tool-initiated events and notifications back to the host. Finally, MCP designs access control and capability negotiation as first-class features, offering a foundation for more auditable and secure AI-to-tool interactions. These architectural differences position MCP as more than just an advanced function-calling mechanism. It shifts the paradigm from tool bindings hardcoded per application toward an interoperable ecosystem of composable, discoverable network services. Since its release, MCP has rapidly grown into a foundational architecture for AI-native applications: thousands of independently developed MCP servers expose model-accessible interfaces to services such as GitHub (Protocol, 2024a), Slack (Protocol, 2024b), and Blender (ahujasid, 2025), while platforms like Cursor (Cursor, 2025) and Claude Desktop (Anthropic, 2024a) demonstrate how MCP-enabled clients can flexibly extend functionality by connecting to new servers on demand. This approach transforms developer tools, productivity platforms, and creative environments into a truly interoperable and multimodal AI ecosystem.
@@
Despite the rapid adoption of MCP, its ecosystem is still in the early stages, with key areas such as security, tool discoverability, and remote deployment lacking comprehensive solutions. These issues present untapped opportunities for further research and development. Although MCP is widely recognized for its potential in the industry, it has not yet been extensively analyzed in academic research. This gap motivates our work, which provides the first in-depth analysis of the MCP ecosystem by examining its architecture and workflow, formally defining the lifecycle of MCP servers across four phases and 16 key activities, and systematically analyzing security threats from multiple attacker perspectives. Our threat taxonomy identifies four major attacker types, including malicious developers, external attackers, malicious users, and security flaws, and covers 16 representative threat scenarios. These risks, such as tool poisoning, installer spoofing, and unauthorized access, are empirically validated through real-world case studies. Through this study, we present a thorough exploration of MCP’s current landscape and offer a forward-looking vision that highlights key implications, outlines future research directions, and addresses the challenges that must be overcome to ensure its sustainable growth.
@@
  • We identify the key components of MCP servers and define their lifecycle, encompassing the stages of creation, deployment, operation, and maintenance, across 16 key activities.
  • We construct the systematic threat taxonomy for MCP, identifying four attacker archetypes (i.e., malicious developers, external attackers, malicious users, and security flaws) and 16 threat scenarios that together reveal the MCP’s primary security exposure points.
@@
The remainder of this paper is structured as follows: § 2 compares tool invocation with and without MCP, highlighting the motivation for this study. § 3 outlines the architecture of MCP, detailing the roles of the MCP host, client, and server, as well as the lifecycle of the MCP server. § 4 examines the current MCP landscape, focusing on key industry players and adoption trends. § 5 analyzes security and privacy risks of the MCP server and proposes mitigation strategies. § 6 explores implications, future challenges, and recommendations to enhance MCP’s scalability and security in dynamic AI environments. § 7 reviews prior work on tool integration and security in LLM applications. Finally, § 8 concludes the whole paper.
@@ 2.1. AI Tooling
Figure 1. Tool invocation with and without MCP. Without MCP, an AI application interacts with external tools and resources such as web services, databases, and local files through specific APIs. With MCP, the AI application functions as an MCP client that communicates with an MCP server using the MCP protocol, which provides a unified interface for tool access.
@@ 2.1.2. Standardized Plugin Interfaces
To reduce the complexity of manual wiring, plugin-based interfaces such as OpenAI ChatGPT Plugins, introduced in November 2023 (OpenAI, 2023a), allowed AI models to connect with external tools through standardized API schemas like OpenAPI. For example, in the OpenAI Plugin ecosystem, plugins like Zapier allowed models to perform predefined actions, such as sending emails or updating CRM records. However, these interactions were often one-directional and could not maintain state or coordinate multiple steps in a task. There are also new LLM app stores (Zhao et al., 2024) that offer a web services plugin store, like ByteDance Coze (ByteDance, 2024) and Tencent Yuanqi (Tencent, 2024). Although these platforms increased the number of tools available, they also produced separate ecosystems where plugins are platform-specific, which restricts limiting cross-platform compatibility and requiring duplicate maintenance efforts. MCP stands out by being open-source and platform-agnostic, supporting bi-directional communication channels that allow not only model-to-tool invocations but also tool-initiated events and notifications, which are not possible in one-directional plugin designs.
@@ 2.1.3. AI Agent Tool Integration
The emergence of AI agent frameworks like LangChain (LangChain, 2022) and similar tool orchestration frameworks provided a structured way for models to invoke external tools through predefined interfaces, improving automation and adaptability (Xi et al., 2023). However, integrating and maintaining these tools remained largely manual, requiring custom implementations and increasing complexity as the number of tools grew. MCP simplifies this process by introducing a protocol-level abstraction that unifies how tools are described and discovered across platforms. This protocol, in contrast to framework-specific integrations, makes it possible for tools created by various developers to work together, eliminating the need for duplicate maintenance and promoting a shared ecosystem.
@@
Contextual information retrieval methods, such as RAG, leverage vector-based search to retrieve relevant knowledge from databases or knowledge bases, enabling models to supplement responses with up-to-date information (Fan et al., 2024; Cuconasu et al., 2024). While this approach addressed the problem of knowledge cutoff and improved model accuracy, it was limited to passive retrieval of information. It did not inherently allow models to perform active operations, such as modifying data. For example, a RAG-based system could retrieve relevant sections from a product documentation database to assist a customer support AI. However, if the AI needed to update customer records or escalate an issue to human support, it can’t take action beyond providing textual responses. MCP extends beyond passive information retrieval by providing a standardized protocol for both retrieval and action, allowing AI systems to combine knowledge access with secure tool execution under a unified framework.
@@ 2.2. Motivation
MCP has rapidly gained traction in the AI community due to its ability to standardize how AI models interact with external tools, fetch data, and execute operations. By addressing the limitations of manual API wiring, plugin interfaces, and agent frameworks, MCP redefines AI-to-tool interactions by enabling interoperable, secure, and maintainable workflows across heterogeneous environments. Unlike prior approaches, MCP incorporates access control, capability negotiation, and schema discovery as protocol primitives, which distinguishes it from conventional function-calling mechanisms. Despite its growing adoption and promising potential, MCP is still in its early stages, with an evolving ecosystem that remains incomplete. Many key aspects, such as security and tool discoverability, are yet to be fully addressed, leaving ample room for future research and improvement. Moreover, while MCP has gained rapid adoption in the industry, it is still largely unexplored in academia.
@@ 3.1. Core Components
Figure 2. The workflow of MCP. A user prompt is processed through a series of stages involving intent analysis, tool selection, and API invocation across the MCP host, client, and server. The MCP server provides tools, resources, and prompts that enable interaction with external data sources such as web services, databases, and local files. The notation “1:1” in the transfer layer indicates a one-to-one communication link between each MCP client and MCP server during request and response exchange.
@@ 3.1.2. MCP Client
In the MCP host–client–server architecture, the MCP client acts as an intermediary component within the host environment, maintaining a one-to-one communication link with its corresponding MCP server. Operating on behalf of the host, the client initiates requests to the MCP server, queries available functions, and retrieves responses that describe the server’s tools, resources, and capabilities. This design enables the host to seamlessly access and utilize external functionalities provided by MCP servers. In addition to managing requests and responses, the MCP client processes notifications from MCP servers, providing real-time updates about task progress and system status. It also performs sampling to gather data on tool usage and performance, enabling optimization and informed decision-making. The MCP client communicates through the transport layer with MCP servers, facilitating secure, reliable data exchange and smooth interaction between the host and external resources.
@@ 3.1.3. MCP Server
  • Tools: Enabling external operations. Tools in MCP enable the server to invoke external services and APIs to execute operations on behalf of AI models. When a host application (such as an AI assistant) needs to perform an operation, it first queries the MCP server through the client to obtain the list of available tools and their capabilities. Based on the task context, the host then selects an appropriate tool and issues an invocation request via the client. The MCP server executes the corresponding operation through the external service or API and returns the result to the client, which forwards it to the host. For example, if an AI model requires real-time weather data or sentiment analysis, the host identifies the corresponding tool from the server’s advertised capabilities, the server performs the API call, and the resulting data is delivered back to the host. Unlike traditional function-calling interfaces that are confined within a single model or framework, tools in MCP are described and accessed through a standardized, model-agnostic protocol. This design allows tools to be dynamically discovered, described, and invoked across heterogeneous AI systems, ensuring cross-platform interoperability rather than provider-specific integration. Furthermore, MCP tools support bi-directional communication between the host and the service, enabling richer interactions such as asynchronous updates or event streaming. Once configured, these tools adhere to a supply-and-consume model that promotes modularity and reusability, allowing independently developed tools to interoperate seamlessly and improving system efficiency and extensibility.
@@ 3.3. MCP Server Lifecycle
Figure 3 summarizes the complete lifecycle of an MCP server from the server’s perspective. This lifecycle is primarily derived from the official protocol documentation and a systematic analysis of actual MCP operational workflows. Based on the transitions of primary participant roles, the lifecycle is divided into four sequential phases: creation, deployment, operation, and maintenance. In the creation phase, the main actor is the developer, who defines metadata, declares capabilities, and implements the MCP server. The deployment phase covers the process in which the developer releases the server to a public platform, and users deploy it to an MCP host while the client establishes its connection. The operation phase corresponds to the runtime period when users actively interact with the server through the MCP system. Finally, the maintenance phase involves version iteration, configuration updates, and continuous optimization of the deployed MCP server. The following subsection introduces the key activities of each phase in detail.
Figure 3. MCP server components and lifecycle. The upper part presents the time-ordered interaction flow among the developer, MCP server, host, client, user, and external resources. The lower part summarizes the main lifecycle phases, including creation, deployment, operation, and maintenance. The numbered circles (e.g., ①②③) indicate sequential actions in the upper process, which correspond to key phases in the lifecycle.
@@ 3.3.1. MCP Server Components
The MCP server is responsible for managing external tools, data sources, and workflows, providing AI models with the necessary resources to perform tasks efficiently and securely. It comprises several key components that ensure smooth and effective operations. Metadata includes essential information about the server, such as its name, version, and description, allowing clients to identify and interact with the appropriate server. Configuration involves the source code, configuration files, and manifest, which define the server’s operational parameters, environment settings, and security policies. Tool list stores a catalog of available tools, detailing their functionalities, input-output formats, and access permissions, ensuring proper tool management and security. Resources list governs access to external data sources, including web APIs, databases, and local files, specifying allowed endpoints and their associated permissions. Finally, Prompt templates include pre-configured task templates and workflows that enhance the efficiency of AI models in executing complex operations. These components enable MCP servers to provide seamless tool integration, data retrieval, and task orchestration for AI-powered applications.
@@ 3.3.2. Creation Phase
The creation phase is the starting point of the MCP server lifecycle. During this stage, developers establish the server’s essential data structures and capabilities, transforming conceptual requirements into a fully defined and executable MCP component. The result is a standardized service that can be reliably integrated into subsequent deployment and operation processes. Metadata definition focuses on describing the server’s identity through its name, version, description, and supported protocol version. These data form the foundation for interoperability and version control. Capability declaration specifies the standardized functions the server provides, such as tools, resources, or prompts, together with their operational boundaries and permission requirements. This step determines how the server will respond to standardized requests. Vague or inaccurate declarations may lead to capability misuse or potential security violations. Many security vulnerabilities identified in MCP deployments originate from flaws introduced at this stage. Therefore, scanning and validation of capability declarations constitute an essential step in strengthening MCP server security. Code implementation connects the declared capabilities with concrete request handlers that process input data, execute corresponding logic, and generate structured results. The quality and security of this implementation directly influence the stability and safety of all exposed functions. Poorly designed or insecure code can introduce vulnerabilities that compromise both reliability and user trust. Slash command definition establishes optional user‑interaction commands that correspond to specific prompts. Clear and coherent command definitions enhance usability and improve the flexibility and intuitiveness of prompt invocation during subsequent operations.
3.3.3. Deployment Phase
In this stage, the server that has been designed and implemented is prepared, packaged, and released into an operational environment where clients and hosts can interact with it through standardized interfaces. MCP server release involves packaging the finalized server codebase, configuration files, and metadata into a distributable form. This step may include version tagging, dependency documentation, and integrity verification to guarantee that the deployed artifact remains consistent with the validated build. Developers can subsequently publish their packaged servers to various MCP server markets, as illustrated in Table 2, allowing end users to discover and install servers directly from these repositories. Most existing markets are maintained by third‑party platforms, but Anthropic has also begun developing an official MCP registry aimed at providing verified listings, enhanced security trust, and unified distribution management for the MCP ecosystem. Installer deployment handles the distribution and installation of the server package within target systems. Installers can use container images, package managers, or automated scripts to streamline deployment. Clear installation workflows reduce setup errors and make the integration process predictable. Security validation of installers, such as checksum verification and signature authentication, is necessary to prevent tampering or injection of malicious binaries. Environment setup ensures that the deployed MCP server operates under the correct runtime configuration. This includes defining environment variables, access credentials, logging policies, and network permissions required for communication with clients and hosts. Configuration isolation and principle‑of‑least‑privilege practices help mitigate risks of unauthorized access or data leakage during runtime. Tool registration finalizes deployment by registering the server’s capabilities with the hosting application or orchestration system. During this process, tools, resources, and prompts declared during creation are discovered and made available to connected clients.